x^2+2x=8 Quadratic Equations
Answer by subudear(62)

x^2+2x=8
x^2+2x-8=0
x^2-2x+4x-8=0
x(x-2)+4(x-2)=0
(x+4)(x-2)=0
therefore, x = 2, -4 is the solution